A 1250 kg car is moving at 12.3 m/s. Then, 325 kJ of work is done on the car to increase its speed. What will be the car's new speed?
Answer #1

WORKDONE = CHANGE IN KINETIC ENERGY

=> W = KEf - KEi

=> KEf = W + KEi = 325*10^3 + [1250*(12.3)^2]/2 = 419556.25 J

=> KEf = (m*v^2)/2 = 419556.25

=> v^2 = 671.29 => v = 25.90 m/s final speed of car.